OpenStudy (anonymous):

You are waiting in line to buy tickets to see your favorite band. You know you will be waiting at least 80 minutes. You are taking some magazines that usually take you 20 minutes apiece to read. a) What inequality symbol would be used to demonstrate this problem? (1 point) b) Write and solve an inequality to determine how many magazines you will need to bring if you plan to read the entire time you wait. (3 points)

OpenStudy (anonymous):

help me plz

OpenStudy (anonymous):

Ok give a sec.....

OpenStudy (anonymous):

me*

OpenStudy (anonymous):

k

OpenStudy (anonymous):

this is so hard and so confusing

OpenStudy (anonymous):

Ok whenever it says 'at least #' Then you know you need to use\[\le or \ge \]For b) each magazine takes 20 min, so you have 20x (x being magazine) and you want to wait the entire time so the equation would look like\[20x \ge 80\] Do you think you can solve from here?

OpenStudy (anonymous):

no

OpenStudy (anonymous):

Ok but do you understand so far?

OpenStudy (anonymous):

kinda not really

OpenStudy (anonymous):

Where do you not understand?

OpenStudy (anonymous):



OpenStudy (anonymous):

everything

OpenStudy (anonymous):

Ok lets try something basic what does this mean\[x \ge 1\]

OpenStudy (anonymous):

its that i just don't know how to do it and put it in a paraph

OpenStudy (anonymous):

i don't know

OpenStudy (anonymous):

Ok so when I was young I would think of the those inequality symbols as an alligator mouth where the alligator would eat the biggest number. So for \[x \ge 1\]You can say x will be eaten by an alligator since its bigger or equal to 1. But in math terms its x is greater than or equal to 1.

OpenStudy (anonymous):

k

OpenStudy (anonymous):

i guess

OpenStudy (anonymous):

So in order to solve\[20x \ge 80\]Just treat the inequality symbol or (alligator mouth) as an equal sign.

OpenStudy (anonymous):

wat!!!!!!!!

OpenStudy (anonymous):

Ok do you know where I got the equation from?

OpenStudy (anonymous):